OpenStudy (anonymous):

factor the sum. 8/27 + m^6

OpenStudy (anonymous):

\[\left( \frac{2}{3}\right)^3 + (m^2)^3 = \left(\frac{2}{3}+m^2\right)\left(\frac{4}{9} + m^4 - \frac{2}{3}\times m^2\right)\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

No Problem